The following matrix is obtained from the augmented matrix of a system of equations.

1 0 3 20
0 1 2 -2
0 0 1 4
What is the solution to the system is?

Answer by stanbon(75887)

1 0 3 20
0 1 2 -2
0 0 1 4
What is the solution to the system is?
----------------
Assuming the variables are a, b, c,
The 3rd line implies that c = 4
If c=4 the 2nd line implies b+2c=-2; b+8=-2; b=-10
If c=4 and b=-10 the 1st line implies a + 3c = 20; a+12=20; a=8
==================
Cheers,
Stan H.